OpenStudy (anonymous):

Calculate the annual wages for a nursing aid who earns $8.29 per hour.

OpenStudy (helder_edwin):

just multiply the hourly wage by the number of hours worked in a year.

OpenStudy (anonymous):

i keep getting 17,246 but it isn't right

OpenStudy (helder_edwin):

say the nurse works 8 hours a day and 5 days a week, it makes 40 hours a week

OpenStudy (anonymous):

uh huh

OpenStudy (helder_edwin):

so in a week she/he makes 40 times $8.29

OpenStudy (helder_edwin):

4 weeks in a month (average) and 12 months then \[ \large $8,29\times40\times4\times12= \]

OpenStudy (anonymous):

$16,580 $15,680 $8,290 $10,120 these are the mc

OpenStudy (anonymous):

but now i have 15916.8

OpenStudy (helder_edwin):

do u know how many hours does a nurse work? anyway it has to be A or B

OpenStudy (anonymous):

\[\left( \frac{ $8.29 }{ 1 hr } \right)\left( \frac{ 8 hr }{ 1day } \right)\left( \frac{ 5day }{ 1week } \right)\left( \frac{ 4week }{ 1month } \right)\left( \frac{ 12month }{ 1yr } \right)=8.29\times8\times5\times4\times12\]

OpenStudy (anonymous):

the unit of answer will be 15916 $ per year if the assumption about how many month does she work in a year is 12 months is correct